OpenStudy (anonymous):

Calculus: How do you differentiate a fraction with a constant in the numerator? Can I use the quotient rule? If so, the constant would be 0, making the entire f' =o, there must be a different method but I can't remember. Help please.

OpenStudy (precal):

 like this

OpenStudy (anonymous):

I guess the asker mean: y = 1/x -> y' = - 1/x²

OpenStudy (anonymous):

Yes, the actual problem is f(x)=24/x^2+12

OpenStudy (anonymous):

so would the constant stay the same?

OpenStudy (precal):

 like this